# Data Mesh: a Systematic Gray Literature Review

### v1

@@ 3.4. Self-Serve Data Platform
Data mesh advocates building domain-agnostic infrastructure and platform services and offering them in a self-service manner to empower different actors involved in creating, deploying, and maintaining data products. By infrastructure services, we meant those related to providing and managing resources such as VMs, networks, and disks. The platform services support design, development, deployment, and management of data products, governance applications, and their components (e.g., data pipelines, ML pipelines, and microservices).
ID4.1. Objectives.
@@
According to the gray literature, the self-serve data platform has several objectives. First, it aims to reduce product developers’ required level of specialization as they do not need to possess in-depth knowledge of managing infrastructure and platform services. Second, it can increase the efficiency of developing and managing products as there is no duplication of efforts from each domain to manage platform services. Last, the common platform services can improve the uniformity of data products, reduce friction between domains, and facilitate interoperability between data products.
ID4.2. Building Self-serve Data Platform.
@@
The platform is built and maintained by a centralized data platform team of highly specialized infrastructure and platform service developers. The central platform team has to be in close contact with all domains to identify and prioritize the requirements for platform services, their features, and their interfaces. The gray literature recommends using IaC (Infrastructure as Code) blueprints to simplify the provision of data products using the self-serve platform. The IaC templates encompass a predefined configuration for infrastructure resources and platform services and integrated security measures, policies, and standards in line with those set by the federated governance team. For example, the data platform team can develop IaC templates for different data product types (e.g., data pipelines, ML pipelines, and ML model services) and different product component types (e.g., a data ingestion component using the Kafka message broker, a data store component using Google object store, and a training pipeline executor component using the AWS SageMaker). The product developers can find the correct IaC templates, customize them as necessary, and use the updated templates to provision and manage their data products.
ID4.3. Components of Self-Serve Data Platform.
From the gray literature, we identified the following key services that self-serve platforms should offer.
ID4.3.1. Computing.:
The platform should offer different computing resources, including hardware accelerators to efficiently execute complex data processing tasks, ML training jobs, and hosting product APIs.
ID4.3.2. Networking.:
The platform should offer various networking capabilities such as virtual private networks, domain name servers, network firewalls, and load balancers.
ID4.3.3. Polyglot Storage.:
@@
Data products might contain different data types that require different data stores. For example, a data product might consist of highly structured, tabular data stored in a relational database. In contrast, other data products might consist of highly unstructured data that can best be stored in an object store (e.g., videos or images). Thus, the platform needs to support storing and serving polyglot data.
ID4.3.4. Services for Data Product Components.:
@@
Data products generally consist of components such as data ingestion, data (ETL) pipeline, ML training pipeline, ML model storage, ML feature storage, message broker, and API. The product developers should be able to use platform services to implement and manage such product components. For example, the developers should be able to use the pipeline orchestration tools from the platform to develop, deploy, schedule, execute, and monitor data and ML pipelines in a self-service manner.
ID4.3.5. Metadata Repository.:
@@
Products, their components, and their artifacts (e.g., ML models and database tables) have various metadata; thus, the platform needs a service to generate, store, and manage metadata. Moreover, company-wide metadata, such as standardized definitions of business terms, should also be available so product developers can incorporate them into their data products.
ID4.3.6. Product/Data Catalog.:
The platform can provide a product catalog service to publish and discover data products. The metadata repository platform service can be used to develop the product catalog.
ID4.3.7. Distributed Query Engine.:
@@
The self-serve platform should offer a distributed query engine that can query data from different domains to enable product aggregators to create composite data products.
ID4.3.8. Monitoring.:
@@
The platform should offer tools that enable domains to monitor their data products and the federated global governance team to monitor the data mesh as a whole. The tools should include but not be limited to functionality that enables monitoring of the operational health, lineage, costs, and performance concerning data quality metrics. They should also support logging, alerts, and automated interventions (e.g., scaling up resources). Finally, the monitoring tools must be integrated with the other relevant platform services, for example, adding monitoring capabilities to data pipelines and ML pipelines.
ID4.3.9. Product Lifecycle Management.:
@@
The platform should offer capabilities (e.g., product versioning, source control, and testing) to allow product developers to manage the lifecycle of a data product. Developers should be able to apply DevOps (Ebert et al., 2016) practices when deploying and operating data products. For example, they should be able to use CI-CD (Continuous Integration and Continuous Delivery or Continuous Deployment) platform services to atomically build and test the data pipelines and APIs used by the products, deploy them using the pipeline orchestration and container hosting services of the self-serve platform.
ID4.3.10. Security and Privacy.:
@@
The platform services must enable realizing data products’ security and privacy requirements. For example, the platform can include a data encryption service to secure data at rest and in motion. In addition, an identity and access management (IAM) service can be provided to manage user and product identities and enforce access control policies for data products and other platform services.
ID4.3.11. Policy Enforcement.:
@@
The platform should offer tools that allow local and global federated governance teams to define, store, attach, observe, and enforce governance policies. The gray literature recommends using the policy-as-code approach, where policies are defined, evaluated, and managed programmatically. Moreover, the platform also needs to provide services to support policy implementation, e.g., data anonymization service and data quality metrics calculation service.
ID4.3.12. Business Intelligence (BI) Tools.:
@@
The platform should enable business users to explore data, create visualizations, generate insights, and create reports in a self-service manner.
Table 6. Benefits, concerns, and applicability: codes and sources.
Sub Category Atomic Codes
ID Name Frequency ID Name Frequency Unique Sources
5.1 Benefits 48 5.1.1 Scalability 12 S3,S4,S12,S13,S18,S25,S28,S48,S60,S62,S69
5.1.2 Increased Agility 7 S3,S25,S28,S40,S92
5.1.3 Higher Data Quality 5 S3,S9,S10,S23,S25
5.1.4 Better Data Discovery 8 S2,S28,S40,S50,S60,S87
5.1.5 Better Governance 6 S3,S10,S12,S23,S28,S95
5.1.6 Reduced Data Lead Time 10 S3,S5,S12,S25,S28,S53,S69,S77
5.2 Concerns 32 5.2.1 Change management 13 S3,S16,S21,S25,S28,S35,S58,S63,S77,S87
5.2.2 Lack of Talent 5 S16,S41,S46,S48,S77
5.2.3 Data Duplication 9 S3,S18,S24,S28,S35,S41,S48,S109
5.2.4 Effort Duplication 5 S4,S9,S18,S35,S85
5.3 Applicability 37 5.3.1 Large and Diverse Data Landscape 20 S3-S5,S7,S9,S10,S12,S22,S25,S26,S58,S62
5.3.2 Need for Agility 7 S3,S4,S5,S12,S22,S35
5.3.3 Need for Better Data Governance 10 S3,S4,S7,S9,S12,S58,S98
4. RQ2. What are the Benefits and Concerns of adopting Data Mesh?
ID5.1. Benefits.
@@ 5.3.3 Need for Better Data Governance 10 S3,S4,S7,S9,S12,S58,S98
From the gray literature, we identified the following potential benefits of data mesh compared to traditional data architectures.
ID5.1.1 Scalability.:
@@
A data mesh architecture is much more scalable than a centralized data architecture since dependencies on a central management platform team are reduced, which allows the organization to scale horizontally (i.e., more data sources) and vertically (i.e., more consumers) without creating a bottleneck.
ID5.1.2. Increased Agility.:
@@
The benefits mentioned above make organizations more agile as they can deliver new data applications at a faster rate and can thus respond quickly to changing needs and goals. Overall, organizations can create more data-driven applications that unlock the value of their data.
ID5.1.3. Higher Data Quality.:
@@
By making those close to the origin of the data responsible for offering it as a data product, accountability for data quality is established. The federated governance team defines a set of global standards and policies that the domains must adhere to when creating data products. Furthermore, domain teams can use their domain-specific knowledge to build data products valuable to various customers.
ID5.1.4. Better Data Discovery:
@@
In the data mesh, domain teams are incentivized to maximize the usage of domain data products and hence to make the products easily discoverable. Therefore, they create self-describing data products and publish product metadata to the central product registry/catalog for easy discoverability.
ID5.1.5. Better Governance.:
@@
Data mesh also allows for better big data governance. The decentralized approach is more suitable for governing a high volume of data by distributing governance responsibilities to the domains close to the origin of the data. As the domains better understand their data, they can better judge who should and should not be able to use certain data in certain conditions and facilitate organizations to comply with external regulations.
ID5.1.6. Reduced Data Lead Time.:
@@
Lastly, domains can independently provision and maintain data products, and consumers can independently use these products through the self-service model, dramatically reducing the data lead time throughout the organization.
ID5.2. Concerns.
@@
Data mesh comes with their own set of challenges. Below, we present an overview of the main concerns for the data mesh recognized by the gray literature.
ID5.2.1. Change management.:
Moving to a data mesh architecture will challenge established organizations from technical and organizational perspectives. Organizations already have established data management processes and a hierarchical organizational structure that will likely need to be restructured.
ID5.2.2. Lack of Talent.:
@@
Data mesh requires domain teams to start operating independently, so the teams need members with the necessary technical skills (e.g., data engineering and machine learning). While a vital goal of the self-serve platform is to empower the developers of less technical expertise to develop data products, building self-serve platform services can be challenging and require time and effort for research and development.
ID5.2.3. Data Duplication.:
@@
As domains repurpose data for new use cases in composite data products, data from source domains may be copied and duplicated. In particular, when implemented on a large scale, data mesh can potentially increase data management costs. Furthermore, as data is stored in different organizational locations, it will become more challenging to govern data.
ID5.2.4. Effort Duplication.:
@@
Even though the self-serve platform tries to reduce the duplication of efforts concerning data platform management, there will still be a certain degree of duplication of efforts and skills for building and maintaining pipelines throughout the different domains compared to a team working on centralized data.
ID5.3. Applicability.
@@
Data mesh may not be the most practical architecture design for all organizations. The gray literature recognizes several conditions for which the data mesh is more suitable than a centralized data architecture.
ID5.3.1. Large and Diverse Data Landscape.:
@@
Organizations with a large data landscape with many data providers and consumers might benefit from a decentralized architecture. Also, when organizations possess diverse data sets that frequently change over time, the data mesh might serve better than a centralized approach with many point-to-point connections.
ID5.3.2. Need for Agility.:
@@
Organizations whose data platform and central data provisioning teams reduce their innovative capabilities and require a faster time to market for data applications can benefit from a data mesh architecture. Also, when there is a need to experiment with data frequently, data mesh is more beneficial than a centralized architecture as it removes the dependency on an oversized and overloaded data provisioning team.
ID5.3.3. Need for better Data Governance.:
@@
Centralized data lakes can quickly become data swamps without data ownership and governance model. Data mesh can resolve these issues for organizations needing clear data ownership and highly governed data based on domain knowledge. In addition, organizations that face many external regulations can use data mesh to ensure compliance as the data mesh facilitates better data governance.

### v2

@@ 3.4. Principle 4: Self-Serve Data Platform
Data mesh advocates building domain-agnostic infrastructure and platform services and offering them in a self-service manner to empower different actors involved in creating, deploying, and maintaining data products (Dehghani, 2022b, 2019b, a). The infrastructure services are to provision and manage computing resources such as VMs, networks, and storage disks. The platform services support the design, development, deployment, and management of data products, governance applications, and their components (e.g., data pipelines, ML pipelines, and microservices).
Objectives.
@@
According to Zhamak (Dehghani, 2022b, 2019b, a), the self-serve platform aims to optimize the experience of data mesh stakeholders (e.g., product developers, consumers, and governance teams) in their respective tasks, i.e., developing, consuming, and governing data products. The gray literature is in agreement with Zhamak. First, the self-serve platform can reduce product developers’ required level of specialization as they do not need in-depth knowledge of managing infrastructure and platform services (Moses, 2020; Furia, 2021; Thandavarayan, 2021; Snowflake, 2021; Baker, 2021a). Second, it can increase the efficiency of developing and managing products as there is no duplication of efforts from each domain to manage platform services (Serra, 2021; Furia, 2021; Snowflake, 2021; Cooke, 2021; Stirling, 2021; Raposo, 2021; Baker, 2021b; Perio, 2022). Last, the common platform services can improve the uniformity of data products, reduce friction between domains, and facilitate interoperability between data products (Lam, 2021; Keboola, 2021; Thoughtworks, 2021; Platter, 2021a).
Building Self-serve Data Platform.
@@
A team of highly specialized infrastructure and platform service developers builds and maintains the platform (Dehghani, 2022a). This central platform team has to be in close contact with all domains to identify and prioritize the requirements for platform services, their features, and their interfaces (Moses, 2022; Kumar, 2022). The gray literature recommends using IaC (Infrastructure as Code) blackprints to simplify the provision of data products using the self-serve platform (Serra, 2021; Nivas Shankar and Hasson, 2021; Cooke, 2021; Strengholt, 2021; Platter, 2021a; Dehghani, 2021a; Ramos, 2021; Maksimavicius, 2021; Wolf, 2022; Strengholt, 2022c). The IaC templates encompass predefined configurations for infrastructure resources and platform services and integrate security measures, policies, and standards in line with those set by the federated governance team. For example, the data platform team can develop IaC templates for different data product types (e.g., data pipelines, ML pipelines, and ML model services) and different product component types (e.g., a data ingestion component using the Kafka message broker, a data store component using Google object store, and a training pipeline executor component using the AWS SageMaker). The product developers can find the correct IaC templates, customize them as necessary, and use the updated templates to provision and manage their data products.
Components of Self-Serve Data Platform.
Zhamak proposes a logical decomposing of the platform into three layers: data infrastructure utility plane, data product experience plane, and mesh experience plane (Dehghani, 2022b, a). They aim to encapsulate the capabilities to manage infrastructure and platform resources, build data products, and govern the data mesh. A few gray literature sources discussed this logical architecture of the self-serve platform (Ramos, 2022). However, many practitioners provided examples for platform components.
Fundamental Computing Resources.:
The platform should offer compute, network, and storage resources, including virtual machines, network devices, and object stores. The literature sources mostly mention using resources from public IaaS (Infrastructure as a Service) providers (Schultze and Wider, 2020b; Nivas Shankar and Hasson, 2021; Snowflake, 2021; Oracle, 2021b; Andrew, 2022; Maharana, 2021; Jochen Christ and Harrer, 2022).
Polyglot Data Stores.:
@@
Regarding data storage, the common recommendation from the practitioners (Jarvin Mutatiina and Blaauw, 2020; Serra, 2021; JPMorgan, 2021; Uttamachandani, 2022; Limited, 2021; Cooke, 2021; Moses, 2021; Wells, 2021; Dehghani, 2021a; Ramos, 2021) and Zhamak (Dehghani, 2019b, 2022a) is polyglot data stores that enable using multiple data storage technologies transparently within a single system. Data products might contain different data types that require different data stores. For example, a data product might consist of highly structured, tabular data stored in a relational database. In contrast, other data products might consist of highly unstructured data that can best be stored in an object store (e.g., videos or images). Thus, the platform must support the storage and serving of polyglot data.
Services for Data Product Components.:
@@
Data products generally consist of components such as data ingestion, data (ETL) pipeline, ML training pipeline, ML model storage, ML feature storage, message broker, and API (Serra, 2021; Devane, 2021; Uttamachandani, 2022; Limited, 2021; Oostra, 2021b; Moses, 2021; Madan, 2021; Baker, 2021a; Raab and Chiapusso, 2021; Ramos, 2021; Wolf, 2022). The product developers should be able to use platform services to implement and manage such product components. For example, the developers should be able to use the pipeline orchestration tools from the platform to develop, deploy, schedule, execute, and monitor data and ML pipelines in a self-service manner.
Metadata Repository/Data Catalog.:
@@
Products, their components, and their artifacts (e.g., ML models and database tables) have various metadata; thus, the platform needs a service to generate, store, and manage metadata. Moreover, company-wide metadata, such as standardized definitions of business terms and reference data, should also be available so product developers can incorporate them into their data products (Strengholt, 2020, 2022a, 2022f). As mentioned in Section 3.1, a data catalog can be used to develop a product catalog service that stores product metadata to enable the discovery of data products.
Federated Query Engine.:
@@
The self-serve platform should offer a federated query engine to query data from different domains and sources, enabling product aggregators to create composite data products easily (Gillin, 2021; Uttamachandani, 2022; Limited, 2021; Ramos, 2021; Jochen Christ and Harrer, 2022).
Monitoring.:
@@
The platform should offer tools that enable domain teams to monitor their data products and the federated global governance team to monitor the data mesh as a whole  (Dehghani, 2022a). The tools should include but not be limited to functionality that enables monitoring of the operational health, lineage, costs, and performance concerning data quality metrics (Moses, 2020; Equation, 2021; Baker, 2021b). They should also support logging, alerts, and automated interventions (e.g., scaling up resources) (Moses, 2020; Keboola, 2021). Finally, the monitoring tools must be integrated with the other relevant platform services, for example, adding monitoring capabilities to data pipelines and ML pipelines (Baker, 2021a).
Product Lifecycle Management.:
@@
The platform should offer capabilities (e.g., product versioning, source control, and testing) to allow product developers to manage the lifecycle of a data product (Moses, 2020; Cooke, 2021; Madan, 2021; Jochen Christ and Harrer, 2022). Developers should be able to apply DevOps (Ebert et al., 2016) practices when deploying and operating data products (Yordanov, 2021; Uttamachandani, 2022). For example, they should be able to use CI-CD (Continuous Integration and Continuous Delivery or Continuous Deployment) platform services to atomically build and test the data pipelines and APIs used by the products, deploy them using the pipeline orchestration and container hosting services of the self-serve platform.
Security and Privacy.:
@@
The platform services must enable the realization of data products’ security and privacy requirements (Dehghani, 2019b, 2022a). For example, the platform can include a data encryption service to secure data at rest and in motion (Nivas Shankar and Hasson, 2021; Perlov, 2021; Kumar, 2022). In addition, an identity and access management (IAM) service can be provided to manage user and product identities and enforce access control policies for data products and other platform services (Serra, 2021; Perlov, 2021; Oracle, 2021a; Reid, 2021; Maharana, 2021; Jochen Christ and Harrer, 2022).
Policy Enforcement.:
@@
The platform should offer tools that allow local and global federated governance teams to define, store, attach, observe, and enforce governance policies (Dehghani, 2019b, 2022a). As mentioned in Section 3.1, the gray literature recommends using the policy-as-code approach for automating policy enforcement. Thus, the platform needs to support necessary tools such as OPA(Open Policy Agent) engines (Oracle, 2021a, b). Moreover, the platform must also provide services to support policy implementation, e.g., data anonymization service and data quality metrics calculation service (Devane, 2021; Orwat, 2021; Raab and Chiapusso, 2021; Dehghani, 2021a; Jochen Christ and Harrer, 2022).
Business Intelligence (BI) Tools.:
@@
BI dashboards and reporting applications are common types of consumer-aligned data products (JPMorgan, 2021; Jochen Christ and Harrer, 2022; Strengholt, 2022b). Thus, the platform should enable business users to explore data, create visualizations, generate insights, and create reports in a self-service manner (Uttamachandani, 2022; Madan, 2021; Adevinta, 2021; Ramos, 2021; Wolf, 2022).
4. RQ2. Motivations, Benefits, Concerns, and Applicability
This section presents the motivation for the data mesh approach, its perceived advantages and disadvantages, and the factors determining its adoption in organizations.
Motivations.
Data Mesh was developed as a direct response to centralized monolithic data management approaches that have been industry standards, such as data warehouses and data lakes. Based on the literature examined in this survey, we find three factors that motivate organizations to adopt a data mesh paradigm in their organization.
Scaling with Data Sources.:
As the number of data sources grows within an organization, it becomes increasingly difficult to maintain an overview of the various sources and guarantee their high-quality delivery. This problem became especially apparent during the rise of big data and has been one of the main motivations for developing data lakes in favor of data warehouses (Giebler et al., 2019). Despite the widespread adoption of data lakes, gray literature demonstrates organizations still struggle to discover and make available all data within the organization  (golombek, 2021; Carr, 2021; Cooke, 2021; Equation, 2021; Tekiner, 2021; Moses, 2021), make various data sources interoperable  (Carr, 2021; Rajagopalan, 2021; Oracle, 2021b; Equation, 2021) and establish data quality and consistency at the source (Jarvin Mutatiina and Blaauw, 2020; Carr, 2021; Thoughtworks, 2021; Rajagopalan, 2021; Thoughtworks, 2020). Data mesh proposes to solve these issues at the source rather than address them centrally.
Central Teams as a Bottleneck.:
The flip side to the ever-increasing number of data sources is the growing number of data-driven use cases in organizations. Usually, the data for these use cases is provided by a centralized team of hyperspecialized data engineers who also manage the monolithic data platforms in which the data is aggregated. However, relying on centralized teams hinders scalability and agility in data delivery, which is required to address the diverse needs of various data consumers (Jarvin Mutatiina and Blaauw, 2020; Roima, 2021; Rajagopalan, 2021; Oostra, 2021b; Snowflake, 2021; Moses, 2020; Rasool, 2021; Bulpin, 2021; Bauer, 2021; Thoughtworks, 2020; Dehghani, 2019a; Thoughtworks, 2021; Schultze and Wider, 2020a; Jochen Christ and Harrer, 2022).
Separation of Provider and Consumer Expertise.:
Another problem, which the previous problem compounds, is that, in most organizations, data provider expertise is separated from data consumer expertise. Effective data usage requires a deep understanding of the data and the proposed use case. Relying on centralized teams to understand both the context of the providing domain and the consuming domain puts additional stress on the existing bottleneck, whereas the data provider and data consumer can be counted on to know their respective domains  (Furia, 2021; Gillin, 2021; Nivas Shankar and Hasson, 2021; Snowflake, 2021).
Benefits.
@@
From the gray literature, we identified the following potential benefits of data mesh compared to traditional data architectures. We briefly discuss each identified potential benefit’s relation to the problems that drive data mesh transition, as identified above.
Scalability.:
@@
A data mesh architecture is much more scalable than a centralized data architecture since dependencies on a central management platform team are reduced, which allows the organization to scale horizontally (i.e., more data sources) and vertically (i.e., more consumers) without creating a bottleneck (Jarvin Mutatiina and Blaauw, 2020; Moses, 2020; Devane, 2021; Nivas Shankar and Hasson, 2021; ELEKS, 2021; Thandavarayan, 2021; Perlov, 2021; Equation, 2021; Bulpin, 2021; one Musa, 2020; Schultze and Wider, 2020b).
Increased Agility.:
@@
The benefits mentioned above make organizations more agile as they reduce the reliance on centralized teams and bring data providers directly in contact with data producers. By removing the bottleneck, they can deliver new data applications at a faster rate and can thus respond quickly to changing needs and goals (Jarvin Mutatiina and Blaauw, 2020; Thandavarayan, 2021; Perlov, 2021; KPMG, 2021; Gaur, 2021). Overall, organizations can create more data-driven applications that unlock the value of their data (Perio, 2022).
Higher Data Quality.:
@@
By making those close to the origin of the data responsible for offering it as a data product, accountability for data quality is established (JPMorgan, 2021; Nivas Shankar and Hasson, 2021; golombek, 2021; Limited, 2021; Bulpin, 2021; Strenholt and Courtright, 2022; Wolf, 2022; Strengholt, 2022d). Domain teams can use their domain-specific knowledge to build data products that are valuable to various customers (Serra, 2021; Carr, 2021; Oostra, 2021b, a). The federated governance team defines a set of global standards and policies that the domains must adhere to when creating data products (Jarvin Mutatiina and Blaauw, 2020; Furia, 2021; Thandavarayan, 2021).
Better Data Discovery:
@@
In the data mesh, domain teams are incentivized to maximize the usage of domain data products and, hence, to make the products easily discoverable (Gaur, 2021; Gleinser, 2021; Bulpin, 2021). Therefore, they create self-describing data products and publish product metadata to the central product registry/catalog for easy discoverability (JPMorgan, 2021; Strenholt and Courtright, 2022). Improving the discoverability of data (products) lets organizations effectively scale the amount of data in their ecosystem.
Better Governance.:
@@
Data mesh also allows for better big data governance. The decentralized approach is more suitable for governing a high volume of data by distributing governance responsibilities to the domains close to the origin of the data (Jarvin Mutatiina and Blaauw, 2020; Devane, 2021; JPMorgan, 2021). Because data providers are in direct contact with data consumers, they can better judge who should and should not be able to use certain data in certain conditions and facilitate organizations to comply with external regulations (Perlov, 2021; Bulpin, 2021; McLarty, 2021).
Reduced Data Lead Time.:
@@
Lastly, domains can independently provision and maintain data products, and consumers can independently use these products through the self-service model without needing to rely on any centralized team. Moreover, domains and federated governance teams are empowered through the self-service platform, dramatically reducing the data lead time throughout the organization (Jarvin Mutatiina and Blaauw, 2020; Cooke, 2021; Oostra, 2021a; Thoughtworks, 2020; Broda, 2022a).
Concerns and Disadvantages.
@@
Data mesh comes with its own set of challenges. Below, we present an overview of the main concerns for the data mesh recognized by the gray literature.
Change management.:
Moving to a data mesh architecture will challenge organizations from technical and organizational perspectives (Jarvin Mutatiina and Blaauw, 2020; Thandavarayan, 2021; Gillin, 2021). Organizations already have established data management processes and a hierarchical organizational structure that will likely need restructured (Gillin, 2021; Perlov, 2021; Rajagopalan, 2021; Strenholt and Courtright, 2022). The employees may be intimidated by changes and resist. To overcome this challenge, the organizations can train their teams on the cultural shift (the way of working) (Rajagopalan, 2021; Shahrin, 2021), educate the stakeholders about the bigger picture and the overall benefits of the transition to the data mesh (Moses, 2021; Shahrin, 2021; Strenholt and Courtright, 2022), and empower the existing business domain teams with data engineers (Shahrin, 2021). The organizations can create an enabling team to work closely with domain teams to help onboard and consume data, train, and guide teams on data mesh best practices (Strengholt, 2022d; Jochen Christ and Harrer, 2022).
Lack of Talent.:
@@
Data mesh requires domain teams to start operating independently, so the teams need members with the necessary technical skills (e.g., data engineering and machine learning). While a vital goal of the self-serve platform is to empower the developers of less technical expertise to develop data products, building self-serve platform services can be challenging and require time and effort for research and development. Unfortunately, there is a lack of experienced data engineers (Gillin, 2021; Saurabh, 2021; Engineering, 2021; Equation, 2021; Thoughtworks, 2020), which was also a common bottleneck in the data warehouse and lake architectures (Thoughtworks, 2020).
Data Duplication.:
@@
As domains repurpose data for new use cases in composite data products, data from source domains may be copied and duplicated. In particular, when implemented on a large scale, data mesh can potentially increase data management costs (Jarvin Mutatiina and Blaauw, 2020; ELEKS, 2021; Hicks, 2021; Thandavarayan, 2021; Perlov, 2021; Strengholt, 2022d). Furthermore, as data is stored in different organizational locations, it will become more challenging to govern data  (Saurabh, 2021; Equation, 2021; Strengholt, 2022d). The remedies proposed by the practitioners include creating data virtualization techniques (e.g., zero-copy data access) that enable ingesting data without copying them and developing open protocols for sharing data securely and efficiently (Rajagopalan, 2021; Saurabh, 2021; Gilmore, 2019).
Effort Duplication.:
@@
Even though the self-serve platform tries to reduce the duplication of efforts concerning data platform management, there will still be a certain degree of duplication of efforts and skills for building and maintaining pipelines throughout the different domains compared to a team working on centralized data (Moses, 2020; Furia, 2021; ELEKS, 2021; Rajagopalan, 2021; Moses, 2022).
Technical Debt.:
Multiple data products and pipelines developed by independent teams may not adhere to a common set of software development best practices and principles (Jarvin Mutatiina and Blaauw, 2020; Riccomini, 2021; Thandavarayan, 2021). Using improper practices when developing data products can lead to a huge technical debt. Hence, a mature federated governance model can reduce the technical debt by accurately recognizing and enforcing standards, best practices, and so on (Jarvin Mutatiina and Blaauw, 2020; Riccomini, 2021; Dehghani, 2022a).
Complexity.:
The distributed and decentralized nature of the data mesh increases the overall complexity as it requires managing and coordinating many independent and diverse units (Oracle, 2021a; Dehghani, 2021a; Strengholt, 2022d). The data mesh can also introduce more interoperability challenges for organizations that adopt a multi-cloud approach (Rajagopalan, 2021). The complexity can decrease as the self-service platform and computational governance model mature (Dehghani, 2022a).
Applicability.
@@
The gray literature recognizes several conditions for which the data mesh is more suitable than a centralized data architecture.
Large and Diverse Data Landscape.:
@@
Organizations with a large data landscape and many data providers and consumers might benefit from a decentralized architecture (Jarvin Mutatiina and Blaauw, 2020; Moses, 2020; Roima, 2021; Serra, 2021; AgileLab, 2021; Mitrus, 2021; Thandavarayan, 2021; Kainos, 2021; Moses, 2021; one Musa, 2020; Lewis, 2021). Also, when organizations possess diverse data sets that frequently change over time, the data mesh might serve better than a centralized approach with many point-to-point connections (Furia, 2021; Mitrus, 2021; Kainos, 2021; Dehghani, 2021a).
Need for Agility.:
@@
Organizations whose data platform and central data provisioning teams reduce their innovative capabilities and require a faster time to market for data applications can benefit from a data mesh architecture (Jarvin Mutatiina and Blaauw, 2020; Moses, 2020; Roima, 2021; Devane, 2021; Mitrus, 2021; Rajagopalan, 2021). Also, when there is a need to experiment with data frequently, data mesh is more beneficial than a centralized architecture as it removes the dependency on an oversized and overloaded data provisioning team (Moses, 2020; Bauer, 2021).
Need for better Data Governance.:
@@
Centralized data lakes can quickly become data swamps without data ownership and governance models. Data mesh can resolve these issues for organizations needing clear data ownership and highly governed data based on domain knowledge (Serra, 2021; Furia, 2021; Moses, 2021). In addition, organizations that face many external regulations can use data mesh to ensure compliance as the data mesh facilitates better data governance (Nivas Shankar and Hasson, 2021; Perlov, 2021; Feingold, 2021).
Data mesh may not be the most practical data architecture for all organizations. The gray literature mentions several scenarios where the data mesh approach may not be the most suitable option.
Lack of Domain-Orientation.:
The data mesh presumes an organizational structure of independent cross-functional teams aligned with the structures of business domains. If an organization (e.g., typically small and medium-sized companies) does not follow domain-driven design principles and does not have multiple teams who can make data products independently, then the data mesh approach may not fit the organization (Roima, 2021; golombek, 2021; Kainos, 2021; Rajagopalan, 2021; Strengholt, 2022d; Jochen Christ and Harrer, 2022).
Low Data Governance Maturity Level.:
The data mesh needs matured data governance support to prevent data silos and to ensure the regulatory compliance of data processing across autonomous domains. Thus, it may not be suitable for organizations with a low level of maturity in data governance (e.g., no automated data quality assessment or regulatory compliance checking) (Devane, 2021; Kainos, 2021; Shahrin, 2021; Strengholt, 2022d; Jochen Christ and Harrer, 2022).
Low Data Platform Maturity Level.:
The self-serve data platform is key to reducing the cost of owning data products and implementing data governance across domains. Building a self-serve data platform can be impractical for organizations without experienced data platform engineers (Kainos, 2021; Jochen Christ and Harrer, 2022).
Lack of Data Use Cases.:
When an organization has a small number of useful analytical data sources and a few data use cases, the cost and risks of transitioning a data mesh would far outweigh the benefits (Devane, 2021; Kainos, 2021; Jochen Christ and Harrer, 2022).